# Bayesian Transfer Operators in Reproducing Kernel Hilbert Spaces

### v1

@@ Appendix D Estimation Algorithm
    This implies that the inverse and inverse square root of the Gramian matrices can be expressed as
@@
D.1 Discussion
Similar to [Baddoo et al., 2022], the Cholesky preconditioning step was also quite beneficial for numerical stability. As a bonus, due to the lower triangular structure of $\mathbf{L}_{ZZ}$ the preconditioned matrices are efficiently found by solving the systems of linear equations.
The jitter term, $\sigma_{Z}^{2}$, implies that we are modeling noisy pseudo-outputs $\bm{f}_{Z}$, and may be regarded as another variational hyperparameter that we can optimize.
As mentioned, TCCA provides another set of quantities by which we can analyze and interpret the dynamical system. The left and right singular functions of the Koopman operator are respectively approximated by $\mathbf{W}_{X}^{\top}\bm{\Psi}_{Z}(\cdot)$ and $\mathbf{W}_{Y}^{\top}\bm{\Psi}_{Z}(\cdot)$, and can be used to identify coherent sets [Froyland et al., 2010, Klus et al., 2019].
@@ D.1 Discussion
As we demonstrated in Section 3, choosing the hyperparameters can be accomplished with a VAMP-score, which is a metric that quantifies the similarity between the estimated singular functions and the true ones. An example is the squared sum of canonical correlations, i.e. the VAMP-$2$ score, which is maximal when all the patterns extracted from the data are perfectly linear, i.e., $\rho_{i}=1\,\forall\,i\,\in\left[\textrm{min}(\textrm{dim}(\mathbf{\Sigma}_{X}),\textrm{dim}(\mathbf{\Sigma}_{Y}))\right]$.

### v2

@@ Appendix D Estimation Algorithm
    This implies that the inverse and inverse square root of the (noisy) Gramian matrices can be expressed as
@@
D.1 Further Remarks
  1. (i)
    Cholesky preconditioning. Similar to [Baddoo et al., 2022], the Cholesky preconditioning step was also quite beneficial for numerical stability. As a bonus, due to the lower triangular structure of $\mathbf{L}_{ZZ}$ the preconditioned matrices are efficiently found by solving the systems of linear equations.
  2. (ii)
    Jitter term and noisy pseudo–outputs. The jitter term, $\sigma_{Z}^{2}$, may be regarded as another variational hyperparameter that we can optimize, and implies that the pseudo-outputs $\bm{f}_{Z}$ are noisy.
  3. (iii)
    TCCA interpretation and Koopman singular functions. As mentioned, TCCA provides another set of quantities by which we can analyze and interpret the dynamical system. The left and right singular functions of the Koopman operator are respectively approximated by $\mathbf{W}_{X}^{\top}\bm{\Psi}_{Z}(\cdot)$ and $\mathbf{W}_{Y}^{\top}\bm{\Psi}_{Z}(\cdot)$, and can be used to identify coherent sets [Froyland et al., 2010, Klus et al., 2019].
@@ D.1 Further Remarks
  4. (iv)
    Hyperparameter selection via VAMP score. As we demonstrated in Section 3, choosing the hyperparameters can be accomplished with a VAMP-score, which is a metric that quantifies the similarity between the estimated singular functions and the true ones. An example is the squared sum of canonical correlations, i.e., the VAMP-$2$ score or the total kinetic variance [Noé and Clementi, 2015], which is a quantity that reaches its maximum when all the patterns extracted from the data are perfectly linear, i.e., $\rho_{i}=1\,\forall\,i\,\in\left[\textrm{min}(\textrm{dim}(\mathbf{\Sigma}_{X}),\textrm{dim}(\mathbf{\Sigma}_{Y}))\right]$.
Appendix E The Bayesian-consistency kernel
Since $\nu_{i}$ is an $\mathbb{H}$-valued Gaussian random element defined by the covariance operator $\mathcal{C}_{\mathrm{bc}}$, i.e.,
$\mathbb{E}\big[\langle\nu_{i},\,g\rangle_{\mathbb{H}}\,\langle\nu_{i},\,f\,\rangle_{\mathbb{H}}\big]=\langle g,\,\mathcal{C}_{\mathrm{bc}}\,f\rangle_{\mathbb{H}}\quad\forall\,f,g\in\mathbb{H}.$
Then the induced scalar process $\nu_{i}(\bm{x})=\langle\nu_{i},\varphi(\bm{x})\rangle_{\mathbb{H}}$ is a Gaussian process associated with the Bayesian-consistency kernel, and can be related to the feature map of the prior covariance function:
$\kappa_{\mathrm{bc}}(\bm{x},\,\bm{x}^{\prime})=\langle\varphi(\bm{x}),\,\mathcal{C}_{\mathrm{bc}}\,\varphi(\bm{x}^{\prime})\rangle_{\mathbb{H}}.$
Since $\mathbf{K}_{\mathrm{bc}}\coloneqq[\kappa_{\mathrm{bc}}(\bm{z}_{i},\bm{z}_{j})]_{i,j=1}^{M}\in\mathbb{R}^{M\times M}$ it follows that
$\mathbf{K}_{\mathrm{bc}}=\Phi_{Z}^{\top}\,\mathcal{C}_{\mathrm{bc}}\,\Phi_{Z}.$
Given a finite-rank approximation of $\mathcal{C}_{\mathrm{bc}}$ supported on the span of $\Phi_{Z}$, such that
$\mathcal{C}_{\mathrm{bc}}\approx\sum_{i,j=1}^{M}\mathbf{C}_{\mathrm{bc},\,ij}\;\varphi(\bm{z}_{i})\otimes\varphi(\bm{z}_{j})=\Phi_{Z}\,\mathbf{C}_{\mathrm{bc}}\,\Phi_{Z}^{\top},\quad\mathbf{C}_{\mathrm{bc}}\in\mathbb{R}^{M\times M},$
we obtain the expression
$\mathbf{C}_{\mathrm{bc}}=\mathbf{K}_{ZZ}^{-1}\,\mathbf{K}_{\mathrm{bc}}\,\mathbf{K}_{ZZ}^{-1}.$